Presented below is the income statement of Goodwin Inc. The income statement is based on sales of 100,000 units at $20 per unit. Goodwin estimates that 80% of Cost of Goods Sold is variable costs, and 80% of Operating expenses is fixed costs. Sales $2,000,000 Less: Cost of goods sold 600,000 Gross margin $1,400,000 Less: Operating expenses 500,000 Net income $ 900,000 Goodwin is considering lowering the sales price in order to increase sales. Management believes that if it reduces the selling price by 10%, then sales (in units) will increase 10%. 8. If Goodwin reduces the selling price by 10% and unit sales increase 10%, net income will: a. Increase b. Decrease 9. Refer to the previous question. What is the amount of the increase or decrease in net income? 10. Assume Goodwin reduces the selling price by 10%. How many units would Goodwin have to sell to earn the same TOTAL contribution margin prior to the reduction in the selling price

Help with Questions 8, 9 & 10
Answer #1
Contribution Margin Income Statement
Current New
Sales        2,000,000        1,980,000
Less: variable cost
Cost of goods sold            480,000            528,000
Operating Expenses            100,000            110,000
Contribution Margin        1,420,000        1,342,000
Less: Fixed costs
Cost of goods sold            120,000            120,000
Operating Expenses            400,000            400,000
Net Income            900,000            822,000
8.b. Decrease
9.Amount of decrease = $78,000
10.Desired Contribution Margin = $1,420,000
Revised Contribution Margin per Unit = 1,342,000/110,000 = $12.2
Hence, units required to be sold = 1,420,000/12.2 = 116,393.44 Units